Courtney measured off a distance of 115 inches. Dylan said, “Okay, I will time how long it takes your car to travel that distance. I have a stop watch feature on my watch.” Courtney raced her remote control car over the 115-inch distance. Dylan said, “Wow, that only took 3.21 seconds. That seems really fast.” “See, I told you,” said Courtney. “But really how fast is it going?” asked Dylan.
Determine the speed of Courtney’s remote control car and compare its speed to that of a regular car traveling on the freeway. Explain how you found your solution.
1
Expert's answer
2020-04-22T19:40:06-0400

Let's determine the speed of Courtney’s remote control car

everyone knows the formula

"\\ V = \frac{S}{t}: V=\frac{115in}{3.21s}=35.8\frac{in}{s}\\"

speed of a regular car can be 50mph

"50 mph\times 17.6 =880\frac{in}{s}"

compare Courtney’s remote control car speed to that of a regular car

"\frac{880}{35.8}=24.58"

"880-35.8=844.2\frac{in}{s}"

Answer: the speed of a regular car is 24.8 times that of a car with a remote control Courtney or

the speed of a normal car is 844.2 in/s faster than the speed of a car with a remote control Courtney